# Token-level Direct Preference Optimization

### v1

@@ 2 Related Works
The emergence of ChatGPT has catalyzed significant advancements in the field of Large Language Models (LLMs), such as OpenAI’s GPT-4 (Achiam et al., 2023), Mistral (Jiang et al., 2023), and Google’s Gemini (Team et al., 2023). Generally, the training of LLMs involves three stages: initial unsupervised pre-training on massive text corpora to grasp linguistic structures (Raffel et al., 2020; Brown et al., 2020; Workshop et al., 2022; Touvron et al., 2023), followed by supervised fine-tuning with task-specific datasets to enhance the LLMs’ probability of producing desired responses (Taori et al., 2023; Chiang et al., 2023; Vu et al., 2023). However, due to the typically limited and expensive availability of labeled datasets during the supervised fine-tuning stage, the model may retain biases and inaccuracies, manifesting as societal biases (Sheng et al., 2021), ethical concerns (Weidinger et al., 2021), toxicity (Rauh et al., 2022), and hallucinations (Huang et al., 2023), which necessitates a subsequent AI alignment phase. Noteworthy models achieving significant alignment, such as Zephyr (Tunstall et al., 2023) and GPT-4 (Achiam et al., 2023), have demonstrated the effectiveness of techniques like RLHF and DPO algorithms.
Reinforcement Learning from Human Feedback (RLHF) has emerged as a cornerstone in aligning LLMs with human values, providing a mechanism to refine model outputs based on qualitative feedback (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022; Song et al., 2023; Touvron et al., 2023). However, the complexity of implementing RLHF, compounded by the inaccuracies in human-generated reward models (Wu et al., 2023), has prompted the exploration of alternative strategies. Methods like Reward Ranked FineTuning (RAFT) (Dong et al., 2023) and Rank Responses to align Human Feedback (RRHF) (Yuan et al., 2023) offer streamlined approaches to alignment, circumventing some of RLHF’s inherent challenges. Particularly, Direct Preference Optimization (DPO) (Rafailov et al., 2023) represents a breakthrough in direct policy optimization, addressing the intricacies of balancing model behavior through a nuanced approach to reward function optimization. Nevertheless, the challenge of maintaining linguistic diversity while aligning with human preferences remains a pivotal concern, prompting our proposed Token-level Direct Preference Optimization (TDPO), which seeks to harmonize the dual objectives of alignment accuracy and expressive range in model outputs.
@@ 4.2 Token-Level Optimization
4.3 Equivalence and Optimization Objective
@@ 5.1 Experiments on IMDb Dataset
The results are depicted in Figure 2(a). We implement the DPO, $\mathrm{TDPO{}}_{1}$, and different versions of $\mathrm{TDPO{}}_{2}$ algorithms with varying the parameter $\alpha$. From the figure, we notice that although DPO establishes an efficient frontier, $\mathrm{TDPO{}}_{1}$ and $\mathrm{TDPO{}}_{2}$ outperform DPO in terms of divergence versus reward on the frontier, achieving higher reward while maintaining low KL divergence. We also implemented versions of $\mathrm{TDPO{}}_{2}$ with $\alpha\in\{1,1.5,2,5\}$. However, we found that higher values of $\alpha$ made it difficult to optimize the reward. In Figures 2(b), 2(c) and 2(d), we illustrate the curves portraying the sequential KL divergence for different algorithms during the training process. The sequential KL divergence growth rate of DPO on the dispreferred response subset is significantly higher than that on the preferred response subset, leading to an increasing offset between them. In contrast, $\mathrm{TDPO{}}_{2}$ exhibits superior control over KL divergence, achieving better divergence efficiency compared to DPO. As analyzed in Section 4.4, $\mathrm{TDPO{}}_{1}$ tends to result in an increased sequential KL divergence on the preferred response subset, thereby exhibiting a weaker capacity for KL divergence adjustment compared to $\mathrm{TDPO{}}_{2}$. $\mathrm{TDPO{}}_{2}$ maintains a more balanced sequential KL divergence on both dispreferred and preferred response subsets, contributing to its ability to achieve a superior frontier. Although a larger $\alpha$ enhances control over the sequential KL divergence, it also affects the speed and difficulty of optimization. For the remainder of this paper, we set $\alpha=0.5$. In Appendix C, we also present graphs of the frontier between the reward and forward KL divergence and the progression curves of the forward KL divergence throughout the training process.
@@ 6 Conclusion
Impact Statements
@@ A.3 Proving the Equivalence of the Bradley-Terry Model and the Regret Preference Model
$P_{\mathrm{BT}}({y}_{1}\succ{y}_{2}|{x})=\sigma\left(\sum_{t=1}^{T_{1}}\gamma^{t-1}A_{\pi}([{x},y_{1}^{<t}],y_{1}^{t})-\sum_{t=1}^{T_{2}}\gamma^{t-1}A_{\pi}([{x},y_{2}^{<t}],y_{2}^{t})\right),$ (34)

### v2

@@ 2 Related Works
The emergence of ChatGPT has catalyzed significant advancements in the field of Large Language Models (LLMs), such as OpenAI’s GPT-4 (Achiam et al., 2023), Mistral (Jiang et al., 2023), and Google’s Gemini (Team et al., 2023). These breakthroughs have pushed the boundaries of what LLMs can achieve, expanding their applications across various domains. Generally, the training of LLMs involves three stages: initial unsupervised pre-training on massive text corpora to grasp linguistic structures (Raffel et al., 2020; Brown et al., 2020; Workshop et al., 2022; Touvron et al., 2023), followed by supervised fine-tuning with task-specific datasets to enhance the LLMs’ probability of producing desired responses (Taori et al., 2023; Chiang et al., 2023; Vu et al., 2023). However, due to the typically limited and expensive availability of labeled datasets during the supervised fine-tuning stage, the model may retain biases and inaccuracies, manifesting as societal biases (Sheng et al., 2021), ethical concerns (Weidinger et al., 2021), toxicity (Rauh et al., 2022), and hallucinations (Huang et al., 2023), which necessitates a subsequent AI alignment phase. Noteworthy models achieving significant alignment, such as Zephyr (Tunstall et al., 2023) and GPT-4 (Achiam et al., 2023), have demonstrated the effectiveness of techniques like Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO) algorithms.
Reinforcement Learning from Human Feedback (RLHF) has emerged as a cornerstone in aligning LLMs with human values, providing a mechanism to refine model outputs based on qualitative feedback (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022; Song et al., 2023; Touvron et al., 2023). This approach has shown considerable promise in making models more responsive to human expectations and ethical considerations by iteratively improving their performance through human-generated feedback. However, the complexity of implementing RLHF, compounded by the inaccuracies in human-generated reward models (Wu et al., 2023), has prompted the exploration of alternative strategies. Methods like Reward Ranked FineTuning (RAFT) (Dong et al., 2023) and Rank Responses to align Human Feedback (RRHF) (Yuan et al., 2023) offer streamlined approaches to alignment, circumventing some of RLHF’s inherent challenges. Particularly, Direct Preference Optimization (DPO) (Rafailov et al., 2023) represents a breakthrough in direct policy optimization, addressing the intricacies of balancing model behavior through a nuanced approach to reward function optimization. Nevertheless, the challenge of maintaining linguistic diversity while aligning with human preferences remains a pivotal concern, prompting our proposed Token-level Direct Preference Optimization (TDPO), which seeks to harmonize the dual objectives of alignment accuracy and expressive range in model outputs.
@@ 4.2 Token-Level Optimization
4.3 BT Model Reformulation via Advantage Function
@@ 4.4 Loss Function and Formal Analysis
where $\alpha$ is a parameter, and
We summarize the comparison of the loss functions for $\mathrm{DPO}$, $\mathrm{TDPO}_{1}$, and $\mathrm{TDPO}_{2}$, as presented in Figure 2.
Figure 2: Comparison of Loss Functions for $\mathrm{DPO}$, $\mathrm{TDPO}_{1}$ and $\mathrm{TDPO}_{2}$ Methods. The $\mathnormal{sg}$ denotes the stop-gradient operator. Both $\mathrm{TDPO}_{1}$ and $\mathrm{TDPO}_{2}$ incorporate an additional term for finer-grained control over the KL divergence, compared to DPO.
@@ 5.1 Experiments on IMDb Dataset
The results are depicted in Figure 3(a). We implement the DPO, $\mathrm{TDPO{}}_{1}$, and different versions of $\mathrm{TDPO{}}_{2}$ algorithms with varying the parameter $\alpha$. From the figure, we notice that although DPO establishes an efficient frontier, $\mathrm{TDPO{}}_{1}$ and $\mathrm{TDPO{}}_{2}$ outperform DPO in terms of divergence versus reward on the frontier, achieving higher reward while maintaining low KL divergence. We also implemented versions of $\mathrm{TDPO{}}_{2}$ with $\alpha\in\{1,1.5,2,5\}$. However, we found that higher values of $\alpha$ made it difficult to optimize the reward. In Figures 3(b), 3(c) and 3(d), we illustrate the curves portraying the sequential KL divergence for different algorithms during the training process. The sequential KL divergence growth rate of DPO on the dispreferred response subset is significantly higher than that on the preferred response subset, leading to an increasing offset between them. In contrast, $\mathrm{TDPO{}}_{2}$ exhibits superior control over KL divergence, achieving better divergence efficiency compared to DPO. As analyzed in Section 4.4, $\mathrm{TDPO{}}_{1}$ tends to result in an increased sequential KL divergence on the preferred response subset, thereby exhibiting a weaker capacity for KL divergence adjustment compared to $\mathrm{TDPO{}}_{2}$. $\mathrm{TDPO{}}_{2}$ maintains a more balanced sequential KL divergence on both dispreferred and preferred response subsets, contributing to its ability to achieve a superior frontier. Although a larger $\alpha$ enhances control over the sequential KL divergence, it also affects the speed and difficulty of optimization. For the remainder of this paper, we set $\alpha=0.5$. In Appendix C, we also present graphs of the frontier between the reward and forward KL divergence and the progression curves of the forward KL divergence throughout the training process.
@@ 6 Conclusion
Regarding the future prospects of alignment methodologies, we anticipate that iterative refinement approaches and multi-turn conversational alignment strategies will significantly improve the alignment of large language models with human values. By continuously refining these models, we can achieve more precise alignment with complex human preferences. Moreover, multi-turn conversations enable deeper and more nuanced interactions, fostering comprehensive attunement to human intentions. These approaches aim to enhance the quality and relevance of AI responses, making AI systems more harmonized with human values and expectations.
Acknowledgements
The research leading to these results received funding from National Key R&D Program of China (2022ZD0116402). In addition, it received funding from Science and Technology Research and Development Project of China State Railway Group Corporation Limited (P2022X012).
Impact Statement
@@ A.3 Proving the Equivalence of the Bradley-Terry Model and the Regret Preference Model
, (34)